I have 1/3 I need to know how many to make 1 gallon?
Answer by Maths68(1474)

Let x is the required amount to add in 1/3 to make 1 gallon
1/3+x=1
subtract 1/3 both sides of the equation
(1/3-1/3)+x=1-1/3
x=1-1/3
taking LCM on right side
x=(3-1)/3
x=2/3 gallon
So we need to add 2/3 gallon to 1/3 gallon to make 1 gallon.
Check
(1/3)+(2/3)=1
Taking LCM on left side
(1+2)/3=1
3/3=1
1=1
You have 1/3 gallon